OpenStudy (anonymous):

x^(5/9) write in radical notation

jhonyy9 (jhonyy9):

\[\sqrt[9]x ^{5}{?}\]

jhonyy9 (jhonyy9):

without ?

OpenStudy (jamesj):

I generally prefer to give hints rather than answers. But this is one of those things you either know or you don't \[ \sqrt[9]{x^5} \] In general (for b > 0), \[ x^{a/b} = \sqrt[b]{x^a} \]

OpenStudy (anonymous):

thankyou